OpenStudy (anonymous):

how would i make this into a single logarithm?? 3logbase(7)x-(1/4)logbase(7)w+4logbase(7)z??

OpenStudy (anonymous):

heres the actuall equation if you dont get my typed ver. \[3\log _{7}x-\frac{ 1 }{ 4 }\log _{7}w+4\log_{7}z\]

OpenStudy (anonymous):

Use the following properties: \[\large a\log_bc=\log_b\left(c^a\right)\\ \large \log_ba+\log_bc=\log_b(ac)\\ \large\log_ba-\log_bc=\log\left(\frac{a}{c}\right)\]

OpenStudy (anonymous):

ok... so i would get \[ \frac{ \log _{7}x ^{3} z ^{4}}{ \log _{7}\sqrt[4]{z} }\]

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

ahhh i see

OpenStudy (anonymous):

that would go to\[\frac{ z ^{\frac{ 19 }{ 4 }} }{ 1 }\]

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

oh ok now that would go outside the fraction right?

OpenStudy (anonymous):

ahhh i see thanks alot :)

OpenStudy (anonymous):

Wait, back up a bit. Mistakes have been made! Where did the \(w\) go?

OpenStudy (anonymous):

\[\log_7\left(\frac{x^3z^4}{w^{1/4}}\right)\] This would be your final answer.

OpenStudy (anonymous):

oh shoot... its the z in the denomitator =0

OpenStudy (anonymous):

ya haha oops... but thanks olot :)

OpenStudy (anonymous):

You're welcome!